# Understanding Stellar Mass-Metallicity and Size Relations in Simulated Ultra-Faint Dwarf Galaxies

### v1

@@
Abstract
Reproducing the physical characteristics of ultra-faint dwarf galaxies (UFDs) in cosmological simulations is challenging, particularly with respect to stellar metallicity and galaxy size. To investigate these difficulties in detail, we conduct high-resolution simulations ($M_{\rm gas}\sim 60~{}\mbox{M}_{\odot}$, $M_{\rm DM}\sim 370~{}\mbox{M}_{\odot}$) on six UFD analogs ($M_{\rm vir}\sim 10^{8}-10^{9}~{}\mbox{M}_{\odot}$, $M_{\rm\star}\sim 10^{3}-2.1\times 10^{4}~{}\mbox{M}_{\odot}$ at $z=0$). Our findings reveal that the stellar properties of UFD analogs are shaped by diverse star-forming environments from multiple progenitor halos in the early Universe. Notably, our UFD analogs exhibit a better match to the observed mass-metallicity relation (MZR), showing higher average metallicity compared to other theoretical models, though our results remain 0.5-1 dex lower than observed UFDs. The metallicity distribution functions (MDFs) of our simulated UFDs lack high-metallicity stars ($\rm[Fe/H]\geq-2.0$) while containing low-metallicity stars ($\rm[Fe/H]<-4.0$) formed primarily through external enrichment. Excluding these low-metallicity stars, our results align well with the MDFs of observed UFDs. However, forming stars with higher metallicity ($\rm-2.0\leq[Fe/H]_{max}\leq-1.5$) remains a challenge due to the difficulty of sustaining metal enrichment during their brief star formation period before cosmic reionization. Additionally, our simulations show extended outer structures in UFDs, similar to recent Milky Way UFD observations, resulting from dry mergers between progenitor halos. To ensure consistency, we adopt the same fitting method commonly used in observations to derive the half-light radius. We find that this method tends to produce lower values compared to direct calculations and struggles to accurately describe the extended outer structures. To address this, we employ a two-component density profile to obtain structural parameters, demonstrating that a two-component exponential profile better describes the galaxy shape, including both inner and outer structures.
@@ 2.1 Simulation Set Up
In this work, we conduct a cosmological hydrodynamic simulation to explore the formation of UFD analogs. We utilize a modified version of parallel N-body/smoothed particle hydrodynamics (SPH) code p-gadget3(Springel et al., 2001; Springel, 2005; Schaye et al., 2010). The initial conditions are generated using the cosmological initial condition generator, music(Hahn & Abel, 2011). To select a target halo at $z=0$, we initially perform a low-resolution DM-only simulation within a box of $L_{\rm box}=3.125\ h^{-1}$ comoving Mpc from $z=128$ to $z=0$. After selecting isolated UFD analogs at $z=0$, we apply a zoom-in technique to the regions encompassing the identified UFD analog. This approach enables us to generate high-resolution initial conditions, achieving a resolution where $M_{\rm{gas}}\approx 60\ {\rm{M_{\odot}}},$ and $M_{\rm{DM}}\approx 370\ {M_{\odot}}$.
@@
We create six initial conditions for isolated UFDs with varying virial masses in the range of $M_{\rm vir}=\rm 10^{8}\sim 10^{9}\ M_{\odot}$. In this setup, the total mass of the most massive halo, Halo6, is approximately ten times that of the least massive halo, Halo1. This mass range enables us to investigate how virial mass influences galactic properties such as stellar mass, average metallicity, SFH, and more. The detailed properties are provided in Table 1. We set the cosmological parameters as follows, a matter density parameter of $\Omega_{\rm m}=1-\Omega_{\Lambda}=0.265$, a baryon density of $\Omega_{\rm b}=0.0448$, a present-day Hubble expansion rate of $\rm H_{\rm 0}=71\,\mathrm{km,s^{-1}Mpc^{-1}}$, a spectral index of $n_{\rm s}=0.963$, and a normalization of $\sigma_{8}=0.8$ (Komatsu et al., 2011; Planck Collaboration, 2016). The adopted softening length in our simulations is $\epsilon_{\rm soft}\sim 20$ physical pc, which remains constant across all redshifts for both DM and baryonic particles. In addition, we use adaptive softening lengths for the gas particles, where the softening length is proportional to the SPH kernel length, with a minimum value of $\epsilon_{\rm gas,min}=2.8\ \rm pc$.
@@
We update the abundances of primordial species ($\rm H,H^{+},H^{-},H_{2},H_{2}^{+},He,He^{+},He^{++},e^{-},D,D^{+}$ and HD) by solving the non-equilibrium rate equations at every time step, incorporating all relevant cooling processes. These processes include H and He collisional ionization, excitation and recombination cooling, bremsstrahlung, inverse Compton cooling, and collisional excitation cooling of $\rm H_{2}$ and HD. Also, we account for metal line cooling involving seven species, C, O, Si, Mg, Ne, N, and Fe. The cooling rates for these metals are interpolated based on density, temperature, and metallicity using tables generated by the photoionization package cloudy(Ferland et al., 1998) (see Wiersma et al. (2009) for more details). Furthermore, we adopt the UV background values from Haardt & Madau (2012), initiating with zero intensity at $z=7$ and linearly increasing it to full strength by $z=6$, under the assumption of complete reionization by $z=6$ (e.g., Gunn & Peterson, 1965; Fan et al., 2006).
@@ 2.2 Star formation
We employ a stochastic star formation method based on the Schmidt law (Schmidt, 1959), where the star formation rate is defined as $\dot{\rho_{\ast}}=\rho/\tau_{\ast}$. In this equation, $\rho$ represents the gas density for star formation, and $\tau_{\ast}$ is given by $\tau_{\rm{ff}}/\epsilon_{\rm{ff}}$, where $\tau_{\rm{ff}}$ is the free-fall time at $\rho$ and $\epsilon_{\rm{ff}}$ is the star formation efficiency. The free-fall time is expressed as $\tau_{\rm ff}=[3\pi/(32G\rho)]^{1/2}$ and $\epsilon_{\rm ff}=0.01$, a value calibrated from observations in the local Universe (e.g., Leroy et al., 2008). Consequently, the star formation timescale can be normalized as follows
@@
When the number density of a gas particle surpasses a threshold of $n_{\rm H,\ thr}=100$ $\rm cm^{-3}$, it transforms into a star particle, retaining the metallicity of the gas. The type of the stellar population is then determined by the metallicity of the gas. We utilize a critical metallicity value of $Z_{\rm crit}=10^{-5.5}Z_{\odot}$. Gas with metallicity above this threshold gives rise to Pop II stars, while gas with $Z_{\ast}\leq Z_{\rm crit}$ leads to the formation of Pop III stars (e.g., Omukai, 2000; Schneider & Omukai, 2010; Safranek-Shrader et al., 2016). If a star particle is identified as a Pop III star, it is treated as an individual star and its mass is sampled from the assumed top-heavy IMF over the mass range $[\rm 1,260]\ M_{\odot}$ according to the following equation,
@@
On the other hand, if a star particle is classified as a Pop II star, we perform individual IMF sampling from a star particle with a mass of $m_{\star}=60M_{\odot}$, thus forming a SSP. We sample masses from the Salpeter IMF (Salpeter, 1955), defined by the equation $dN/d\log m\approx m^{-\alpha}$, with a slope $\alpha=1.35$ over the mass range $[0.1-100]M_{\odot}$, until the total mass reaches $60M_{\odot}$. If, during this sampling, a mass falls within the range of $8~{}\mbox{M}_{\odot}-40M_{\odot}$, the eligible range for a Type II SN explosion, we designate this particle as a Type II SN progenitor with the sampled mass. This approach of individual star sampling, where a single stellar particle can result in a discrete SN explosion, allows us to accurately capture the discrete nature of SN explosions. If the star particle does not contain a mass within the eligible range for an SN explosion, it consists of low-mass stars totaling $60M_{\odot}$ and does not trigger any SN explosion.
@@ 2.3 Stellar feedback
Since we treat Type II SN explosions individually for both Pop II and Pop III stars, metal enrichment by Type II SN is also achieved on an individual basis. For Pop III stars that end their lives as CCSNe, the amount of ejected mass and metal yields are provided by Heger & Woosley (2010). For those that die as PISNe, the values are given by Heger & Woosley (2002). Pop II stars yield their metals through three processes: AGB stars, CCSNe, and Type Ia SNe. When CCSNe occur, the metals are ejected based on values from Portinari et al. (1998). Stars with masses in the range of $0.8\ M_{\odot}-8M_{\odot}$ undergo AGB mass loss, with yields provided by Marigo (2001) as a function of their stellar age, metallicity, and mass. Type Ia SNe yields are determined based on the delay time distribution for stars within the mass range of $3\ M_{\odot}-8\ M_{\odot}$. Given the uncertainties in the detailed evolution of Type Ia SN, we adopt an empirical delay-time function characterized by e-folding times (Förster et al., 2006). Notably, AGB and Type Ia SN yields are not ejected from individual stars but from a SSP star with a total mass of 60 $\rm M_{\odot}$, thus yielding IMF-averaged values at a given time. The expelled metals are initially transported among neighboring gas particles, approximately $N_{\rm Ngb}\approx 48$, and then diffused onto the ambient medium using the method described by Greif et al. (2009). This diffusion scheme determines the degree of mixing on a sub-resolution scale based on physical properties such as velocity dispersion and density within the SPH kernel of gas particles.
@@ 3 Results
3.1 Morphology and mass growth history
In this section, our analysis primarily focuses on the early epoch at $z\approx 6$. During this period, the progenitor halos that will eventually merge into the UFD analog by $z=0$ remain distinct. This separation allows us to examine how the environment of each progenitor halos affects the stellar properties formed within it. Also, this epoch is appropriate because our target UFDs had already finished their star formation by this time. We identify the progenitor halos at $z\approx 6$ using the halo-finder code rockstar (Behroozi et al., 2013). To determine which progenitor halos will eventually merge by $z=0$, we compare the IDs of DM particles in each progenitor halo at $z\approx 6$ with those of particles within a UFD analog at $z=0$. A progenitor halo is classified as a to-be-accreted progenitor if more than 90% of its DM particles are found within the halo at $z=0$. This results in the number of contributing progenitors being 1, 5, 4, 5, 6, and 3 for each UFD analog from Halo1 to Halo6. These progenitors are selected based on their contribution of more than 5% to the total stellar mass.
Fig. 1 illustrates the DM projection along the z-direction at $z\approx 6$, showing identified progenitor halos distributed at local density peaks. Different length scales are applied to each set for optimal illustration. The progenitor halos are depicted as circles with radii corresponding to their virial radii and are color-coded for distinction. The progenitors are numbered according to their virial mass at $z\approx 6$, with the most massive progenitor in each set labeled as prog1 (red). Progenitors with a stellar mass contribution below 5% of the total stellar mass are uniformly colored grey in the plot. We find that more massive halos tend to have more star-forming progenitors. For example, as shown in Fig. 1, Halo1 has only one star-forming progenitor, prog1. The lack of star formation in other progenitors might be due to Halo1’s relatively small virial mass of $M_{\rm vir,\ z=0}\approx 1.8\times 10^{8}~{}\mbox{M}_{\odot}$, which results in slower mass growth compared to other halos. In contrast, Halo5, which is approximately ten times more massive than Halo1 with a mass of $M_{\rm vir,\ z=0}\approx 1.17\times 10^{9}~{}\mbox{M}_{\odot}$, has a total of 14 star-forming progenitor halos, including progenitor halos with stellar mass less than 5% of the total. However, this trend is not strictly linear, as the most massive Halo6, showing a greater virial mass than Halo5 at $M_{\rm vir,\ z=0}\approx 1.27\times 10^{9}~{}\mbox{M}_{\odot}$, has 7 star-forming progenitors, which is half the number of progenitors in Halo5.
@@ 3.1 Morphology and mass growth history
Fig. 2 shows the evolution of the virial mass (solid lines) and stellar mass (dashed lines) of progenitor halos, using the same color scheme as in Fig. 1. When calculating the stellar mass, we exclude stars that have exploded as SNe, since these stars lose their mass in the explosion and do not significantly contribute to the final stellar mass. Our results indicate that progenitor halos begin their star formation at different times, with more massive halos tending to form stars earlier. The variations in the timing of star formation are primarily driven by differences in the mass growth histories of the progenitor halos, as a progenitor halo needs to exceed a virial mass of $M_{\rm vir}=10^{5}-10^{6}~{}\mbox{M}_{\odot}$ for gas clouds to collapse and form stars (e.g., Bromm & Larson, 2004; Hirano et al., 2015). For instance, the earliest star formation occurs at $z\sim 16$ in prog3 of Halo4, where the virial mass reaches $M_{\rm vir}\sim 6.5\times 10^{6}~{}\mbox{M}_{\odot}$, the highest mass among the progenitor halos in Halo4 at this epoch. Here, a Pop III star forms first, followed by a Pop II star at $z\sim 15.7$, after a delay of 6 Myr (e.g., Ritter et al., 2012; Jeon et al., 2014). On the other hand, prog1 of Halo1 begins star formation later at $z\sim 7.8$, with a virial mass of $M_{\rm vir}\sim 6.3\times 10^{6}~{}\mbox{M}_{\odot}$. The viral masses of progenitor halos at the time of first star formation range from $M_{\rm vir}\approx 1.1\times 10^{6}~{}\mbox{M}_{\odot}$ to $6.6\times 10^{6}~{}\mbox{M}_{\odot}$, reflecting the varying environments of each progenitor halo. Furthermore, it is important to note that the first formed star is not always a Pop III star; it could be a Pop II star if the interstellar medium (ISM) of the progenitor halo is externally enriched. In this work, 90% of progenitor halos have a Pop III star as their first formed star, with only 4 out of 40 progenitor halos experiencing external enrichment at the time of their first star formation.
@@
We find that star formation in the progenitor halo does not occur continuously over long periods but rather in episodic bursts. During these episodic phases, the stellar mass increases rapidly and then pauses until the next burst of star formation begins. This episodic nature is attributed to the shallow potential wells of the progenitor halos, which are unable to sustain star formation for extended periods due to stellar feedback. During a starburst, stellar feedback heats the surrounding ISM and expels gas from the central region of the progenitor halo, temporarily suppressing or even halting further star formation. This delayed star formation is particularly pronounced in prog2 of Halo4, where star formation begins at $z\sim 10$ and resumes at $z\sim 7$ after a delay of about 300 Myr. All the simulated UFD analogs complete their star formation by $z\sim 6$, as the ISM is heated above $\rm T\approx 10^{4}$ K once the UV background is introduced at $z\sim 7$, in combination with the local SN feedback effect.
@@
Determining the main progenitor halo, when defined as the most massive progenitor at a given time, proves to be challenging because it changes over time. Fig. 3 shows the stellar mass of each progenitor halo, with numbering and colors matching those in Fig.  1. A prominent feature in Fig. 3 is that the main progenitor (prog1) at $z\approx 6$ does not significantly contribute to the total stellar mass. For Halo2 to Halo5, the stellar mass fractions of the main progenitor at $z=6$ are 31.1%, 52.3%, 45.7%, 15.0%, and 48.0%, respectively. This suggests that in the simulated UFD analogs, stars formed in the main progenitor account for less than 50% of the total stellar mass. Furthermore, in the case of Halo5, the most substantial contributor to the stellar mass is prog5, which contributes 29% of the total stellar mass of Halo5. This implies that caution is needed when assuming that most stars in the observed UFD originate from a single progenitor, especially when deducing their SFHs or stellar environments.

### v2

@@
Understanding Stellar Mass-Metallicity and Size Relations in Simulated Ultra-Faint Dwarf Galaxies
Minsung Go School of Space Research, Kyung Hee University, 1732 Deogyeong-daero, Yongin-si, Gyeonggi-do 17104, Republic of Korea Myoungwon Jeon School of Space Research, Kyung Hee University, 1732 Deogyeong-daero, Yongin-si, Gyeonggi-do 17104, Republic of Korea Department of Astronomy & Space Science, Kyung Hee University, 1732 Deogyeong-daero, Yongin-si, Gyeonggi-do 17104, Republic of Korea Myoungwon Jeon myjeon@khu.ac.kr Yumi Choi NSF National Optical-Infrared Astronomy Research Laboratory, 950 North Cherry Avenue, Tucson, AZ 85719, USA Nitya Kallivayalil Department of Astronomy, The University of Virginia, 530 McCormick Road, Charlottesville, VA 22904, USA Sangmo Tony Sohn Space Telescope Science Institute, 3700 San Martin Drive, Baltimore, MD 21218, USA Department of Astronomy & Space Science, Kyung Hee University, 1732 Deogyeong-daero, Yongin-si, Gyeonggi-do 17104, Republic of Korea Gurtina Besla Department of Astronomy, University of Arizona, 933 N. Cherry Ave, Tucson, AZ 85721, USA Hannah Richstein Department of Astronomy, The University of Virginia, 530 McCormick Road, Charlottesville, VA 22904, USA Sal Wanying Fu Department of Astronomy, University of California, Berkeley, Berkeley, CA, 94720, USA Tae Bong Jeong School of Space Research, Kyung Hee University, 1732 Deogyeong-daero, Yongin-si, Gyeonggi-do 17104, Republic of Korea Jihye Shin Korea Astronomy and Space Science Institute, 776 Daedeokdae-ro, Yuseong-gu, Daejeon 34055, Republic of Korea
(Received 21 Nov 2024; Revised 15 Apr 2025; Accepted 29 Apr 2025)
@@ 2.1 Simulation Set Up
In this work, we conduct a cosmological hydrodynamic simulation to explore the formation of UFD analogs. We utilize a modified version of parallel N-body/smoothed particle hydrodynamics (SPH) code p-gadget3 (Springel et al., 2001; Springel, 2005; Schaye et al., 2010). The initial conditions are generated using the cosmological initial condition generator, music (Hahn & Abel, 2011). To select a target halo at $z=0$, we initially perform a low-resolution DM-only simulation within a box of $L_{\rm box}=3.125\ h^{-1}$ comoving Mpc from $z=128$ to $z=0$. After identifying isolated UFD analogs at $z=0$, we apply a zoom-in technique to the region surrounding the identified UFD analog. This region is determined by particles within $3\ R_{\rm vir}$ at $z=0$, ensuring that boundary effects in the Lagrangian region do not impact the properties of our target UFD analogs. As a result, we generate high-resolution initial conditions, achieving a resolution where $M_{\rm{gas}}\approx 60\ {\rm{M_{\odot}}},$ and $M_{\rm{DM}}\approx 300\ ~{}\mbox{M}_{\odot}$.
@@
We create six initial conditions for isolated UFDs with varying virial masses in the range of $M_{\rm vir}=\rm 10^{8}\sim 10^{9}\ M_{\odot}$. In this setup, the total mass of the most massive halo, Halo6, is approximately ten times that of the least massive halo, Halo1. This mass range enables us to investigate how virial mass influences galactic properties such as stellar mass, average metallicity, SFH, and more. The detailed properties are provided in Table 1. We set the cosmological parameters as follows, a matter density parameter of $\Omega_{\rm m}=1-\Omega_{\Lambda}=0.265$, a baryon density of $\Omega_{\rm b}=0.0448$, a present-day Hubble expansion rate of $\rm H_{\rm 0}=71\,\mathrm{km,s^{-1}Mpc^{-1}}$, a spectral index of $n_{\rm s}=0.963$, and a normalization of $\sigma_{8}=0.8$ (Komatsu et al., 2011; Planck Collaboration, 2016). The adopted softening length in our simulations is $\epsilon_{\rm soft}\sim 20$ physical pc, which remains constant across all redshifts for both DM and star particles. In addition, we use adaptive softening lengths for the gas particles, where the softening length is proportional to the SPH kernel length, with a minimum value of $\epsilon_{\rm gas,min}=2.8\ \rm pc$.
@@
We update the abundances of primordial species ($\rm H,H^{+},H^{-},H_{2},H_{2}^{+},He,He^{+},He^{++},e^{-},D,D^{+}$ and HD) by solving the non-equilibrium rate equations at every time step, incorporating all relevant cooling processes. These processes include collisional ionization of H and He, excitation and recombination cooling, bremsstrahlung, inverse Compton cooling, and collisional excitation cooling of $\rm H_{2}$ and HD. In addition, we account for metal line cooling involving seven species, C, O, Si, Mg, Ne, N, and Fe. The cooling rates for these metals are interpolated based on density, temperature, and metallicity using tables generated by the photoionization package cloudy (Ferland et al., 1998) (see Wiersma et al. (2009) for more details). Furthermore, we adopt the UV background values from Haardt & Madau (2012), initiating with zero intensity at $z=7$ and linearly increasing it to full strength by $z=6$, under the assumption of complete reionization by $z=6$ (e.g., Gunn & Peterson, 1965; Fan et al., 2006). Self-shielding of dense gas is implemented by attenuating the UV background according to $\exp(-N_{\rm HI}\ \overline{\sigma}_{\rm ion})$, where the neutral hydrogen column density is given by $N_{\rm H\ I}={\rm x}\ n_{\rm HI}$. Here, $\rm x=h$ represents the SPH kernel size, $n_{\rm HI}$ is the number density of neutral hydrogen, and $\overline{\sigma}_{\rm ion}$ denotes the frequency-averaged photoionization cross section for $\rm HI$. For more details, we refer readers to Jeon et al. (2017)
@@ 2.2 Star formation
We employ a stochastic star formation method based on the Schmidt law (Schmidt, 1959), where the star formation rate (SFR) is defined as $\dot{\rho_{\ast}}=\rho/\tau_{\ast}$. In this equation, $\rho$ represents the gas density for star formation, and $\tau_{\ast}$ is given by $\tau_{\rm{ff}}/\epsilon_{\rm{ff}}$, where $\tau_{\rm{ff}}$ is the free-fall time at $\rho$ and $\epsilon_{\rm{ff}}$ is the star formation efficiency. The free-fall time is expressed as $\tau_{\rm ff}=[3\pi/(32G\rho)]^{1/2}$ and $\epsilon_{\rm ff}=0.01$, a value calibrated from observations in the local Universe (e.g., Leroy et al., 2008). Consequently, the star formation timescale can be normalized as follows
@@
When the number density of a gas particle surpasses a threshold of $n_{\rm H,\ th}=100$ $\rm cm^{-3}$, it transforms into a star particle, retaining the metallicity of the gas. To assess the dependence of stellar mass on different parameters, $\epsilon_{\rm ff}$ and $n_{\rm H,th}$, we used the same initial conditions for Halo4 and performed two additional test runs with $\epsilon_{\rm ff}=0.3$ and $\epsilon_{\rm ff}=1.0$, respectively. Furthermore, we performed another run with $n_{\rm H,th}=500~{}\mathrm{cm^{-3}}$, keeping $\epsilon_{\rm ff}$ fixed at 0.01 to isolate the effect of the increased density threshold. The resulting stellar masses varied by a factor of two compared to the default setup ($\epsilon_{\rm ff}=0.01,\ n_{\rm H,th}=100~{}\mathrm{cm^{-3}}$) and no clear trend is observed—for example, increasing $\epsilon_{\rm ff}$ did not lead to a corresponding increase in stellar mass. Previous studies have shown that the stellar mass is insensitive to the adopted star formation parameters (e.g., Hopkins et al., 2014; Gutcke et al., 2021), and similar insensitivity might be expected in UFD analogs. However, due to their extremely low stellar masses, UFD analogs are more significantly affected by stochastic processes, which likely hinders the convergence of stellar mass across different runs.
The type of stellar population is then determined by the metallicity of the gas. We utilize a critical metallicity value of $Z_{\rm crit}=10^{-5.5}Z_{\odot}$. Gas with metallicity above this threshold gives rise to Pop II stars, while gas with $Z_{\ast}\leq Z_{\rm crit}$ leads to the formation of Pop III stars (e.g., Omukai, 2000; Schneider & Omukai, 2010; Safranek-Shrader et al., 2016). If a star particle is identified as a Pop III star, it is treated as an individual star and its mass is sampled from the assumed top-heavy IMF over the mass range $[\rm 1,260]\ M_{\odot}$ according to the following equation,
@@
On the other hand, if a star particle is classified as a Pop II star, we perform individual IMF sampling from a star particle with a mass of $m_{\star}=60~{}\mbox{M}_{\odot}$, thus forming a SSP. We sample masses from the Salpeter IMF (Salpeter, 1955), defined by the equation $dN/d\log m\approx m^{-\alpha}$, with a slope $\alpha=1.35$ over the mass range $[0.1-100]~{}\mbox{M}_{\odot}$, until the total mass reaches $60~{}\mbox{M}_{\odot}$. If, during this sampling, a mass falls within the range of $[8-40]~{}\mbox{M}_{\odot}$, the eligible range for a Type II SN explosion, we designate this particle as a Type II SN progenitor with the sampled mass.
Since the individually sampled star particles have masses within $8~{}\mbox{M}_{\odot}-100~{}\mbox{M}_{\odot}$, we use the following method to conserve the original gas-particle mass of $60~{}\mbox{M}_{\odot}$. If a sampled star has a mass $m_{\star}\in[8-60]~{}\mbox{M}_{\odot}$, we assign a lost mass of $60-m_{\star}$ to the next SSP particle. Conversely, if a star with $m_{\star}\in[60,100]~{}\mbox{M}_{\odot}$, we assign a reduced mass of $120-m_{\star}$ to the next SSP particle. This approach of individual star sampling, where a single stellar particle can result in a discrete SN explosion, allows us to accurately capture the discrete nature of SN explosions. If the star particle does not contain a mass within the eligible range for a SN explosion, it consists of low-mass stars totaling $60~{}\mbox{M}_{\odot}$ and does not trigger any SN explosion.
Revaz et al. (2016) highlighted that incomplete sampling of the IMF using high-resolution SSP particles can increase scatter in alpha abundance ratios, especially for low-metallicity stars, and as well as MDF. However, our study shows that star formation primarily occurs in starbursts over short timescales, minimizing the impact of Type Ia SNe and AGB star losses on stellar metallicity since our SSP is a 60 ${}\mbox{M}_{\odot}$ particle with stars ranging from 0.1 to 8 ${}\mbox{M}_{\odot}$, and the tendency for stars to form in groups helps average out the effects of incomplete IMF sampling.
@@ 2.3 Stellar feedback
Since we treat Type II SN explosions individually for both Pop II and Pop III stars, metal enrichment by Type II SN is also achieved on an individual basis. For Pop III stars that end their lives as CCSNe, the amount of ejected mass and metal yields are provided by Heger & Woosley (2010). For those that die as PISNe, the values are given by Heger & Woosley (2002). Pop II stars yield their metals through three processes: AGB stars, CCSNe, and Type Ia SNe. When CCSNe occur, the metals are ejected based on values from Portinari et al. (1998). Stars with masses in the range of $0.8\ ~{}\mbox{M}_{\odot}-8~{}\mbox{M}_{\odot}$ undergo AGB mass loss, with yields provided by Marigo (2001) as a function of their stellar age, metallicity, and mass. Type Ia SNe yields are determined based on the delay time distribution for stars within the mass range of $3\ ~{}\mbox{M}_{\odot}-8\ ~{}\mbox{M}_{\odot}$. Given the uncertainties in the detailed evolution of Type Ia SN, we adopt an empirical delay-time function characterized by e-folding times (Förster et al., 2006). Notably, AGB and Type Ia SN yields are not ejected from individual stars but from a SSP star with a total mass of 60 $\rm M_{\odot}$, thus yielding IMF-averaged values at a given time. The expelled metals are initially transported among neighboring gas particles, approximately $N_{\rm Ngb}\approx 48$, and then diffused onto the ambient medium using the method described by Greif et al. (2009). This diffusion scheme determines the degree of mixing on a sub-resolution scale based on physical properties such as velocity dispersion and density within the SPH kernel of gas particles.
@@ 3 Results
3.1 Morphology and mass growth history
In this section, our analysis primarily focuses on the early epoch at $z\approx 6$. During this period, the progenitor halos that will eventually merge into the UFD analog by $z=0$ remain distinct. Due to the physical separation of progenitor halos from one another, we can infer that the stellar properties in each progenitor halo are influenced solely by its own environment, without influence from other progenitor halos. Also, this epoch is appropriate because our target UFDs had already finished their star formation before $z\approx 6$. We identify the progenitor halos at this epoch using the halo-finder code rockstar (Behroozi et al., 2013). To determine which progenitor halos will eventually merge by $z=0$, we compare the IDs of DM particles in each progenitor halo at $z\approx 6$ with those of particles within a UFD analog at $z=0$. A progenitor halo is classified as a to-be-accreted progenitor if more than 90% of its DM particles are found within the halo at $z=0$. This results in the number of contributing progenitors being 1, 5, 4, 5, 6, and 3 for each UFD analog from Halo1 to Halo6. These progenitors are selected based on their contribution of more than 5% to the total stellar mass.
Fig. 1 illustrates the DM projection along the z-direction at $z\approx 6$, showing identified progenitor halos distributed at local density peaks. Different length scales are applied to each set for optimal illustration. The progenitor halos are depicted as circles with radii corresponding to their virial radii and are color-coded for distinction. The progenitors are numbered according to their virial mass at $z\approx 6$, with the most massive progenitor in each set labeled as prog1 (red). Progenitors with a stellar mass contribution below 5% of the total stellar mass are uniformly colored grey in the plot.
@@ 3.1 Morphology and mass growth history
We find that the number of star-forming progenitors is associated with the final halo mass at $z=0$. For example, as shown in Fig. 1, Halo1 has only one star-forming progenitor halo, prog1, likely due to Halo1’s relatively small virial mass of $M_{\rm vir,\ z=0}\approx 1.8\times 10^{8}~{}\mbox{M}_{\odot}$. In contrast, Halo5, which is approximately ten times more massive than Halo1 with a mass of $M_{\rm vir,\ z=0}\approx 1.17\times 10^{9}~{}\mbox{M}_{\odot}$, has a total of 14 star-forming progenitor halos, including those with stellar mass less than 5% of the total. However, this trend is not strictly linear, as the most massive Halo6, showing a greater virial mass than Halo5 at $M_{\rm vir,\ z=0}\approx 1.27\times 10^{9}~{}\mbox{M}_{\odot}$, has 7 star-forming progenitors, which is half the number of progenitors in Halo5.
@@
Fig. 2 shows the evolution of the virial mass (solid lines) and stellar mass (dashed lines) of progenitor halos, using the same color scheme as in Fig. 1. When calculating the stellar mass, we exclude stars that have exploded as SNe, since these stars lose their mass in the explosion and do not significantly contribute to the final stellar mass. Our results indicate that progenitor halos begin their star formation at different times, with more massive halos tending to form stars earlier. The variations in the timing of star formation are primarily driven by differences in the mass growth histories of the progenitor halos, as a progenitor halo needs to exceed a virial mass of $M_{\rm vir}=10^{5}-10^{6}~{}\mbox{M}_{\odot}$ for gas clouds to collapse and form stars (e.g., Bromm & Larson, 2004; Hirano et al., 2015). For instance, the earliest star formation occurs at $z\sim 16$ in prog3 of Halo4, where the virial mass reaches $M_{\rm vir}\sim 6.5\times 10^{6}~{}\mbox{M}_{\odot}$, the highest mass among the progenitor halos in Halo4 at this epoch. Here, a Pop III star forms first, followed by a Pop II star at $z\sim 15.7$, after a delay of 6 Myr (e.g., Ritter et al., 2012; Jeon et al., 2014). On the other hand, prog1 of Halo1 begins star formation later at $z\sim 7.8$, with $M_{\rm vir}\sim 6.3\times 10^{6}~{}\mbox{M}_{\odot}$. The viral masses of progenitor halos at the time of first star formation range from $M_{\rm vir}\approx 1.1\times 10^{6}~{}\mbox{M}_{\odot}$ to $6.6\times 10^{6}~{}\mbox{M}_{\odot}$, reflecting the varying environment of each progenitor halo. Furthermore, it is important to note that the first-formed star is not always a Pop III star; it could be a Pop II star if the interstellar medium (ISM) within the progenitor halo is externally enriched. In this work, 90% of progenitor halos have a Pop III star as their first formed star, with only 4 out of 40 progenitor halos experiencing external enrichment at the time of their first star formation.
@@
We find that star formation in the progenitor halo does not occur continuously over long periods but rather in episodic bursts. During these episodic phases, the stellar mass increases rapidly and then pauses until the next burst of star formation begins. This episodic nature is attributed to the shallow potential wells of the progenitor halos, which are unable to sustain star formation for extended periods due to stellar feedback. During a starburst, stellar feedback heats the surrounding ISM and expels gas from the central region of the progenitor halo, temporarily suppressing or even halting further star formation. This delayed star formation is particularly pronounced in prog2 of Halo4, where star formation begins at $z\sim 10$ and resumes at $z\sim 7$ after a delay of about 300 Myr.
@@
Fig. 3 shows the stellar masses of each progenitor halos, with numbering and colors matching those in Fig. 1. Here, a prominent feature is that the main progenitor [^2] We define a main progenitor as the most massive one at a given time. Note that, in some cases, the main progenitor halo changes over time according to their mass growth histories. (prog1) at $z\approx 6$ does not significantly contribute to the total stellar mass. For Halo2 to Halo5, the stellar mass fractions of the main progenitor at $z=6$ are 31.1%, 52.3%, 45.7%, 15.0%, and 48.0%, respectively. This suggests that in the simulated UFD analogs, stars formed in the main progenitor account for less than 50% of the total stellar mass. Furthermore, in the case of Halo5, the most substantial contributor to the stellar mass is prog5, which contributes 29% of the total stellar mass of Halo5. This implies that caution is needed when assuming that most stars in the observed UFD originate from a single progenitor, especially when deducing their SFHs or stellar environments.